# fastkqr: A Fast Algorithm for Kernel Quantile Regression

### v1

@@ E.6 Proof of Theorem 7
$\left\{\begin{array}[]{l}\mathbf{0}=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i},\\ 0=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t},\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},i\in\hat{S}_{t},\end{array}\right.$ (57)
@@
$\left\{\begin{array}[]{l}\mathbf{0}=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ 0=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t},\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},i\in\hat{S}_{t}.\end{array}\right.$ (58)
@@
$\begin{split}\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}&\sum^{T}_{t=1}\left[\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}})+\frac{\lambda_{2}}{2}\boldsymbol{\alpha}_{\tau_{t}}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]\\ &+\lambda_{1}\sum^{T-1}_{t=1}V(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}})\\ \text{subject to}\enskip{}&y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,i\in\hat{S}_{t},t=1,\ldots,T.\end{split}$ (59)
We see that $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ satisfies the KKT conditions of the convex problem (59). Therefore, $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ is the minimizer of problem (59). Moreover, it can be easily seen that $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}$ is a feasible point of (59) since $\hat{S}_{t}\subseteq{}S_{0}$. This implies that

### v2

@@ 7 Disclosure Statement
$\left\{\begin{array}[]{l}\mathbf{0}=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i},\\ 0=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t},\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},i\in\hat{S}_{t},\end{array}\right.$ (S59)
@@
$\left\{\begin{array}[]{l}\mathbf{0}=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ 0=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t},\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},i\in\hat{S}_{t}.\end{array}\right.$ (S60)
@@
$\begin{split}\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}&\sum^{T}_{t=1}\left[\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}})+\frac{\lambda_{2}}{2}\boldsymbol{\alpha}_{\tau_{t}}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]\\ &+\lambda_{1}\sum^{T-1}_{t=1}V(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}})\\ \text{subject to}\enskip{}&y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,i\in\hat{S}_{t},t=1,\ldots,T.\end{split}$ (S61)
We see that $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ satisfies the KKT conditions of the convex problem (S61). Therefore, $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ is the minimizer of problem (S61). Moreover, it can be easily seen that $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}$ is a feasible point of (S61) since $\hat{S}_{t}\subseteq{}S_{0}$. This implies that